OpenStudy (anonymous):

14. A cartoon plane with four engines can accelerate at 8.9 m/s2 when one engine is running. What is the acceleration of the plane if all four engines are running and each produces the same force?

OpenStudy (azureilai):

F=ma if one engine produce 8.9m/s2 of acceleration, how do you think you can figure out 4 engines (especially if they have the same force)?

OpenStudy (anonymous):

F=F1+F2+F3+F4=mA1+mA2+mA3+mA4=m(a1+a2+a3+a4),F=4(a1),4(8.9)=35.6?

OpenStudy (azureilai):

yep

OpenStudy (anonymous):

Is it okay if you can help me on:The surface gravity of the sun is 274 m/s^2.How many Earth g's is this? If not I can close this question ^^

ganeshie8 (ganeshie8):

Earth g = 9.8

ganeshie8 (ganeshie8):

divide

OpenStudy (anonymous):

Thanks!

ganeshie8 (ganeshie8):

np :) wat did u get when u divide ?

OpenStudy (anonymous):

27.96g's

ganeshie8 (ganeshie8):

\(\large \color{red}{\checkmark}\)